# ATI: Any Trajectory Instruction for Controllable Video Generation

### v2

@@ 1 Introduction
Recent advances in video generation models  [3, 10, 13, 16, 1, 22, 44, 23, 17, 2, 4, 27, 37, 33] have demonstrated remarkable capabilities in synthesizing realistic and diverse video content. However, precise control over motion remains a significant challenge, particularly when users require fine-grained direction over how specific elements move within the generated sequence. Current approaches  [20, 41, 31, 30, 25, 35] typically address different types of motion control—such as camera movement, object translation, or local deformation—through separate specialized modules, leading to fragmented workflows and inconsistent results. Motion control in video generation encompasses a spectrum of manipulations, from camera operations (panning, zooming, rotation) to object-level translations and localized movements of specific regions. These motion types are inherently related and often need to be coordinated to achieve desired visual effects. The separation of these controls in existing systems limits creative expression and requires users to navigate multiple interfaces or models.

### v3

@@ 1 Introduction
Recent advances in video generation models  [3, 11, 14, 17, 1, 23, 44, 24, 18, 2, 4, 28, 37, 34, 9] have demonstrated remarkable capabilities in synthesizing realistic and diverse video content. However, precise control over motion remains a significant challenge, particularly when users require fine-grained direction over how specific elements move within the generated sequence. Current approaches  [21, 41, 32, 31, 26] typically address different types of motion control—such as camera movement, object translation, or local deformation—through separate specialized modules, leading to fragmented workflows and inconsistent results. Motion control in video generation encompasses a spectrum of manipulations, from camera operations (panning, zooming, rotation) to object-level translations and localized movements of specific regions. These motion types are inherently related and often need to be coordinated to achieve desired visual effects. The separation of these controls in existing systems limits creative expression and requires users to navigate multiple interfaces or models.